# Double strangeness pentaquarks and other exotic hadrons in the $\Xi_{b}\to\Xi\ J/\Psi\ \phi$ decay

### v1

@@ VI Conclusions
Finally, we also studied the $\phi~{}\Xi$ spectrum, where we explored the possibility to detect the $\Xi(2500)$. For this resonance there exists a discussion about whether its width is $50~{}{\rm MeV}$ or $150~{}{\rm MeV}$ PDG , so we studied these two possibilities plus an intermediate case of $\Gamma=100~{}{\rm MeV}$. Our results show that, if the resonance has a narrow width of $50~{}{\rm MeV}$, its signal in the spectrum has high chances to be detected even with small coupling $(g_{\phi\Xi}\sim 0.7-0.9)$, whereas if $\Gamma=100~{}{\rm MeV}$ the coupling should be bigger than $g_{\phi\Xi}=1.3$, and in the case $\Gamma=150~{}{\rm MeV}$ the value of the coupling has to be bigger than $g_{\phi\Xi}=1.9$ in order to produce a significant signal above the background.
@@ Acknowledgments
This work is supported by MICIU/AEI/10.13039/ 501100011033 and by FEDER UE through grants PID2023-147112NB-C21; and through the “Unit of Excellence María de Maeztu 2020-2023" award to the Institute of Cosmos Sciences, grant CEX2019-000918-M. Additional support is provided by the Generalitat de Catalunya (AGAUR) through grant 2021SGR01095.
@@ Appendix A Appendix A: Double loop integral
Figure 30: Feynman diagram associated to the ${\cal M}_{4160}^{J/\psi\Xi}$ amplitude, where we show the corresponding momenta for all the particle lines.
In Fig. 30 we can see the the diagram associated to the ${\cal M}_{4160}^{J/\psi\Xi}$, where the momenta of each particle is shown. From this diagram we can write the full expression of the ${\cal M}_{4160}^{J/\psi\Xi}$ amplitude, being:
@@
where we have defined $K_{1}=(P_{\Xi}+P_{\phi})/2$ and $K_{2}=(P_{\Xi}-P_{\phi})/2$ (and correspondingly $K_{1}+K_{2}=P_{\Xi}$ and $K_{1}-K_{2}=P_{\phi}$). $T_{J/\psi\Xi,J/\psi\Xi}$ denotes the scattering amplitude for the corresponding channels, which we introduce as a Breit-Wigner with the $S=-2$ pentaquark parameters. As it is discussed in section IV.3 the $J/\psi$ $\phi$ interaction has to be in P-wave what leads to appearance of the vector product of the corresponding polarization vectors. Finally, the three propagators are the ones corresponding to the three internal lines for the $J/\psi$, $\Xi$ and $X(4160)$ resonances. From this expression we can separate the $q^{0}$ component and rearrange some terms in order simplify it, obtaining that
@@
where $\omega_{X}^{2}=M_{X}^{2}+(\vec{K}_{1}-\vec{q})^{2}$, $\omega_{J/\psi}^{2}=M_{J/\psi}^{2}+(\vec{K}_{2}-\vec{q})^{2}$ and $\omega_{\Xi}^{2}=M_{\Xi}^{2}+(\vec{K}_{2}+\vec{q})^{2}$. If we take a non relativistic limit this expression can be written like this,
@@
$\int\frac{dq^{0}}{2\pi}\frac{G_{D_{s}^{*}\bar{D}_{s}^{*}}}{2\omega_{X}[K_{1}^{0}-q^{0}-\omega_{X}+i\frac{\Gamma_{X}}{2}}$
$\frac{1}{2\omega_{J/\psi}[K_{2}^{0}-q^{0}-\omega_{J/\psi}+i\epsilon]}\frac{M_{\Xi}}{\omega_{\Xi}[K_{2}^{0}+q^{0}-\omega_{\Xi}+i\epsilon]}.$
From this expression we can now clearly see where are the poles of this integral, which are
@@
$q^{0}=K_{2}^{0}-\omega_{J/\psi}+i\epsilon$$\rightarrow$ ${\rm J/\psi~{}propagator},$ (47)
$q^{0}=-K_{2}^{0}+\omega_{\Xi}-i\epsilon$$\rightarrow$ ${\rm\Xi~{}propagator}.$ (48)
Since we have two poles above the real axis and one below, we compute the integral using the contour integration method choosing to close the contour in the lower half plane. Thus we only pick the $\Xi$ propagator pole, finding
@@
$\int\frac{d^{3}q}{(2\pi)^{3}}\frac{(\vec{K}_{2}+\vec{q})G_{D_{s}^{*}\bar{D}_{s}^{*}}}{2\omega_{X}[K_{1}^{0}+K_{2}^{0}-\omega_{\Xi}-\omega_{X}+i\frac{\Gamma_{X}}{2}}$
@@
To compute this integral we work in the Jackson frame (the $J/\psi$ $\phi$ rest frame), where the $\vec{K}_{1}$ and $\vec{K}_{2}$ are almost equal since, $P_{\phi}$ is small. So we can can replace $\vec{K}_{1}$ with $\vec{K}_{2}$, and thus now the integral only depends on $\vec{K}_{2}$. So we can take the integral to be proportional to this vector, giving us
@@
where we made the shift $\vec{q}\to\vec{q}+\vec{K}_{2}$, and $\omega_{X}=M_{X}^{2}+(2\vec{K}_{2}-\vec{q})^{2}$, $\omega_{J/\psi}=M_{J/\psi}^{2}+(2\vec{K}_{2}-\vec{q})^{2}$ and $\omega_{\Xi}=M_{\Xi}^{2}+\vec{q}^{2}$.
Once we have ${\cal M}_{4160}^{J/\psi\Xi}$, to evaluate the ${\cal M}_{4160}^{\phi\Xi}$ is rather straightforward, as we only need to change $J/\psi$ to $\phi$ and vice versa. In the Jackson frame this is equal to the interchange of $\vec{K}_{2}$ and $\vec{K}_{1}$, leading us to
@@
Finally we can also derive the expressions for ${\cal M}_{4140}^{J/\psi\Xi}$, and ${\cal M}_{4140}^{\phi\Xi}$ proceeding in an analogous way with the difference that now we do not have the $G_{D^{*}_{s}\bar{D}^{*}_{s}}$ loop-function and all interactions are in S-wave, and consequently these amplitudes are scalars.
@@ Appendix B Appendix B: Spin Sums
In order to calculate the final mass distribution of the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay we needed to compute the average $\overline{|{\cal M}_{4160}}|^{2}$. In this appendix we show in more details how we arrived to our final result, eq. (IV.4). So we start with the ${\cal M}_{4160}$ amplitude in the following form:
@@
$(P_{\Xi}^{i}\tilde{{\cal M}}_{4160}^{P}+K_{2}^{i}\tilde{{\cal M}}_{4160}^{J/\psi}+K_{1}^{i}\tilde{{\cal M}}_{4160}^{\phi})\cdot$
$\cdot(P_{\Xi}^{a}\tilde{{\cal M}}_{4160}^{*P}+K_{2}^{a}\tilde{{\cal M}}_{4160}^{*J/\psi}+K_{1}^{a}\tilde{{\cal M}}_{4160}^{*\phi}),$ (59)
@@
Note that, since we are working in a reference frame where the $J/\psi~{}\phi$ system is in rest, the 3-momenta of all these particles are small in comparison with their masses, therefore we can take a non-relativistic limit, where
@@
Hence, performing the sum over the polarization, Eq. (B) can be simplified to:
$\overline{|{\cal M}_{4160}|^{2}}=\varepsilon^{ijk}\varepsilon^{ajk}(P_{\Xi}^{i}\tilde{{\cal M}}_{4160}^{P}+K_{2}^{i}\tilde{{\cal M}}_{4160}^{J/\psi}+K_{1}^{i}\tilde{{\cal M}}_{4160}^{\phi})\cdot$
$\cdot(P_{\Xi}^{a}\tilde{{\cal M}}_{4160}^{*P}+K_{2}^{a}\tilde{{\cal M}}_{4160}^{*J/\psi}+K_{1}^{a}\tilde{{\cal M}}_{4160}^{*\phi}).$
@@
$\overline{|{\cal M}_{4160}|^{2}}=(\vec{P}_{\Xi}\tilde{{\cal M}}_{4160}^{P}+\vec{K}_{2}\tilde{{\cal M}}_{4160}^{J/\psi}+\vec{K}_{1}\tilde{{\cal M}}_{4160}^{\phi})\cdot$
$(\vec{P}_{\Xi}\tilde{{\cal M}}_{4160}^{*P}+\vec{K}_{2}\tilde{{\cal M}}_{4160}^{*J/\psi}+\vec{K}_{1}\tilde{{\cal M}}_{4160}^{*\phi})$
. (62)
From this expression we can straightforward derive Eq. (36) just by computing the scalar products.

### v2

@@ VI Conclusions
Finally, we also studied the $\phi~{}\Xi$ spectrum, where we explored the possibility to detect the $\Xi(2500)$. For this resonance there exists a discussion about whether its width is $50~{}{\rm MeV}$ or $150~{}{\rm MeV}$ [7], so we studied these two possibilities plus an intermediate case of $\Gamma=100~{}{\rm MeV}$. Our results show that, if the resonance has a narrow width of $50~{}{\rm MeV}$, its signal in the spectrum has high chances to be detected even with small coupling values of $g_{\phi\Xi}\sim 0.7-0.9$, whereas if $\Gamma=100~{}{\rm MeV}$ the coupling should be larger than $g_{\phi\Xi}=1.3$, and even larger ($g_{\phi\Xi}\gtrsim 1.9$) if $\Gamma=150~{}{\rm MeV}$, in order to produce a significant signal above the background.
@@ Acknowledgments
This work is supported by the State Research Agency of the Ministry of Science, Innovation and Universities of Spain (MICIU/AEI/10.13039/ 501100011033) and by FEDER UE through grant PID2023-147112NB-C21 and through the “Unit of Excellence María de Maeztu 2020-2023" award to the Institute of Cosmos Sciences, grant CEX2019-000918-M. Additional support is provided by the Generalitat de Catalunya (AGAUR) through grant 2021SGR01095.
@@ Appendix A Appendix A: Double loop integral
Figure 29: Feynman diagram associated to the ${\cal M}_{4160}^{J/\psi\Xi}$ amplitude, showing the momenta for all particle lines.
In Fig. 29 we can see the diagram associated to the amplitude ${\cal M}_{4160}^{J/\psi\Xi}$, where the momenta of each particle is shown. The corresponding expression is:
@@
where we have defined $K_{1}=(P_{\Xi}+P_{\phi})/2$ and $K_{2}=(P_{\Xi}-P_{\phi})/2$ (and correspondingly $K_{1}+K_{2}=P_{\Xi}$ and $K_{1}-K_{2}=P_{\phi}$). The $J/\psi\Xi$ scattering amplitude $T_{J/\psi\Xi,J/\psi\Xi}$ is introduced as a Breit-Wigner with the $S=-2$ pentaquark parameters. As discussed in section IV.3 the weak interaction has to be in P-wave what leads to the vector product of the corresponding polarization vectors times the momentum of the $\Xi$. Finally, the three internal lines correspond to the propagators of the $J/\psi$, $\Xi$ and $X(4160)$ resonances. Separating the $q^{0}$ integral and rearranging some terms we obtain:
@@
where $\omega_{X}^{2}=M_{X}^{2}+(\vec{K}_{1}-\vec{q})^{2}$, $\omega_{J/\psi}^{2}=M_{J/\psi}^{2}+(\vec{K}_{2}-\vec{q})^{2}$ and $\omega_{\Xi}^{2}=M_{\Xi}^{2}+(\vec{K}_{2}+\vec{q})^{2}$. Taking a non relativistic limit, this expression can be written as:
@@
$\int\frac{dq^{0}}{2\pi}\frac{G_{D_{s}^{*}\bar{D}_{s}^{*}}}{2\omega_{X}[K_{1}^{0}-q^{0}-\omega_{X}+i\frac{\Gamma_{X}}{2}]}$
$\frac{1}{2\omega_{J/\psi}[K_{2}^{0}-q^{0}-\omega_{J/\psi}+i\epsilon]}\frac{M_{\Xi}}{\omega_{\Xi}[K_{2}^{0}+q^{0}-\omega_{\Xi}+i\epsilon]}\ ,$
where the poles of the integrand can be easily located:
@@
$q^{0}=K_{2}^{0}-\omega_{J/\psi}+i\epsilon$$\rightarrow$ $J/\psi~{}{\rm propagator},$ (46)
$q^{0}=-K_{2}^{0}+\omega_{\Xi}-i\epsilon$$\rightarrow$ $\Xi~{}{\rm propagator}.$ (47)
Since there are two poles above the real axis and one below, we compute the integral closing the contour in the lower half plane, thus picking only the $\Xi$ propagator pole and finding:
@@
$\int\frac{d^{3}q}{(2\pi)^{3}}\frac{(\vec{K}_{2}+\vec{q})G_{D_{s}^{*}\bar{D}_{s}^{*}}}{2\omega_{X}[K_{1}^{0}+K_{2}^{0}-\omega_{\Xi}-\omega_{X}+i\frac{\Gamma_{X}}{2}]}$
@@
To compute this integral we work in the Jackson frame (the $J/\psi$ $\phi$ rest frame), where $\vec{K}_{1}$ and $\vec{K}_{2}$ are almost equal since $P_{\phi}$ is small. So we can can replace $\vec{K}_{1}$ with $\vec{K}_{2}$, and thus now the integral only depends on $\vec{K}_{2}$. In this way the integral can be taken to be proportional to this vector, giving:
@@
where the shift $\vec{q}\to\vec{q}+\vec{K}_{2}$ has been made, and $\omega_{X}=M_{X}^{2}+(2\vec{K}_{2}-\vec{q})^{2}$, $\omega_{J/\psi}=M_{J/\psi}^{2}+(2\vec{K}_{2}-\vec{q})^{2}$ and $\omega_{\Xi}=M_{\Xi}^{2}+\vec{q}^{2}$.
Once we have ${\cal M}_{4160}^{J/\psi\Xi}$, the evaluation of ${\cal M}_{4160}^{\phi\Xi}$ is rather straightforward, as one only needs to change $J/\psi$ to $\phi$ and vice versa. In the Jackson frame this is equal to the interchange of $\vec{K}_{2}$ and $\vec{K}_{1}$, leading to:
@@
Finally, we can also derive the expressions for ${\cal M}_{4140}^{J/\psi\Xi}$ and ${\cal M}_{4140}^{\phi\Xi}$ proceeding in an analogous way, with the difference that now we do not have the $G_{D^{*}_{s}\bar{D}^{*}_{s}}$ loop-function and all interactions are in S-wave, and consequently these amplitudes are scalars. Hence, we can write:
@@ Appendix B Appendix B: Spin Sums
In order to calculate the final mass distribution of the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay we needed to compute the average $\overline{|{\cal M}_{4160}}|^{2}$. In this appendix we show in more detail how we arrived to our final result, Eq. (IV.4). We start with the ${\cal M}_{4160}$ amplitude in the following form:
@@
$\times$ $(P_{\Xi}^{i}\tilde{{\cal M}}_{4160}^{P}+K_{2}^{i}\tilde{{\cal M}}_{4160}^{J/\psi}+K_{1}^{i}\tilde{{\cal M}}_{4160}^{\phi})$
$\times$ $(P_{\Xi}^{a}\tilde{{\cal M}}_{4160}^{*P}+K_{2}^{a}\tilde{{\cal M}}_{4160}^{*J/\psi}+K_{1}^{a}\tilde{{\cal M}}_{4160}^{*\phi}),$
@@
Note that, since we are working in a reference frame where the $J/\psi~{}\phi$ system is at rest, the 3-momenta of all these particles are small in comparison with their masses, therefore we can take a non-relativistic limit, where
@@
Hence, performing the sum over the polarizations, Eq. (B) can be simplified to:
$\overline{|{\cal M}_{4160}|^{2}}$ $=$ $\varepsilon^{ijk}\varepsilon^{ajk}(P_{\Xi}^{i}\tilde{{\cal M}}_{4160}^{P}+K_{2}^{i}\tilde{{\cal M}}_{4160}^{J/\psi}+K_{1}^{i}\tilde{{\cal M}}_{4160}^{\phi})$ (60)
$\times$ $(P_{\Xi}^{a}\tilde{{\cal M}}_{4160}^{*P}+K_{2}^{a}\tilde{{\cal M}}_{4160}^{*J/\psi}+K_{1}^{a}\tilde{{\cal M}}_{4160}^{*\phi}).$
@@
$\overline{|{\cal M}_{4160}|^{2}}=(\vec{P}_{\Xi}\tilde{{\cal M}}_{4160}^{P}+\vec{K}_{2}\tilde{{\cal M}}_{4160}^{J/\psi}+\vec{K}_{1}\tilde{{\cal M}}_{4160}^{\phi})$
×$\times(\vec{P}_{\Xi}\tilde{{\cal M}}_{4160}^{*P}+\vec{K}_{2}\tilde{{\cal M}}_{4160}^{*J/\psi}+\vec{K}_{1}\tilde{{\cal M}}_{4160}^{*\phi}).$ (61)
From this expression we can straightforwardly derive Eq. (35) by just computing the scalar products.